OpenStudy (i<3mybaby):

The angle of elevation of a ladder leaning against a wall is 48°. If the ladder is 25 feet long, how far is the base of the ladder from the wall? A. 18.58 feet B. 16.73 feet C. 37.16 feet D. 33.46 feet @Squirrels @mathmate @mathmale @jcpd910 @dan815

OpenStudy (dan815):



OpenStudy (dan815):

USE SOH CAH TOA

OpenStudy (dan815):

COS 48 deg = A/25 solve for A

OpenStudy (i<3mybaby):

Sin opposite hyp cos adjacent hyp tan opp adj

OpenStudy (i<3mybaby):

for the cos 48 i got 0.6691360635885821382627333068678

OpenStudy (anonymous):

Cos(48) = Adjacent/Hypotenuse = Base/25

OpenStudy (anonymous):

If you want to solve for base you must do 1 more step

OpenStudy (anonymous):

0.6991306 = Base/25

OpenStudy (anonymous):

This is you're result so far

OpenStudy (i<3mybaby):

it is 16.73 I had to have the angle against the wall (45) and the length of the hypotenuse (25ft) and you want to find the adjacent side. so for this we need to use 'cos' (CAH = cos(angle) =adjacent/hypotenuse) therefore: cos(45)=?/25 rearrange to get 25cos(45)=? type into calculator to get 16.73 feet

OpenStudy (anonymous):

Correct

OpenStudy (i<3mybaby):

Jumps up and down YAYAYA